OpenStudy (anonymous):

Diane can type 100 letters in 6hrsm with frankie they can complete the job in 4hrs how long would it take frankie alone?

OpenStudy (anonymous):

Im getting 100/6+100/x=1/4 as the equation not sure if correct

OpenStudy (anonymous):

Got the answer 12, can someone confirm

OpenStudy (anonymous):

will it not be100/6+100/x=100/4

OpenStudy (anonymous):

12 is correct

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw